OpenStudy (anonymous):

simplify and reduce expression to lowest terms 15z/10z+13 - 5z-13/10z+13

OpenStudy (anonymous):

@Ceann

OpenStudy (anonymous):

This is a nice and easy one. Starting with \[\frac{15z}{10z+13}-\frac{5z-13}{10z+13}\] We just do standard fraction subtraction, and it simplifies nicely. \[\frac{10z+13}{10z+13}=1\]